OpenStudy (anonymous):

x^(-1/2)=1/2 how can you get the x alone?

OpenStudy (anonymous):

\[\frac{1}{\sqrt{x}}=\frac{1}{2}\iff \sqrt{x}=2\iff x=4\]

OpenStudy (anonymous):

knowing that \(x^{-r}=\frac{1}{x^r}\) and that \(x^{\frac{1}{2}}=\sqrt{x}\)

OpenStudy (anonymous):

i understand. how do you get the 2 in the second step?

OpenStudy (anonymous):

do you multipy by two? and if so would that not take away the square root?

OpenStudy (anonymous):



Parth (parthkohli):

Or the logarithm trick would work...

OpenStudy (anonymous):